OpenStudy (meilendurcer):

A scuba diver is descending at a rate of 15 3/4 feet per minute. What is the diver’s change in altitude after 4 minutes? A. 60 ft B. -60 ft. C. 63 ft. D. -63 ft.

OpenStudy (anonymous):

You need to multiply 15 3/4 feet per minute by 4 minutes to find the answer.

OpenStudy (anonymous):

You can do multiplying mixed fractions a couple of ways: 1) convert 15 3/4 to an improper fraction like: 15 = 60/4, so 15 3/4 = 63/4 Then multiply 63/4 times 4 minutes 2) multiply 4 x 15 and add 4 x (3/4)

OpenStudy (anonymous):

Alright, the 4 x 15 isn't difficult. That's 60. The easiest way for me to get 3/4 of anything, is to put it in decimal form, and multiply your number by that. 3/4 can be expressed in decimal form as .75, so if you can't see what 3/4 of 4 is in your head, then just multiply 4 by .75 in a calculator. Then add it to 60, and there's your answer. (: